### v1

@@ A.3 Flexible Job-shop Scheduling Problem
Disjunctive Graph for FJSP. To simplify the disjunctive graph for FJSP, we introduce operation-machine nodes, where an operation is decomposed into multiple operation-machine nodes. Each operation-machine node is similar to a node in JSP and can be treated as an action. Specifically, the disjunctive graph is formalized as $G=(V,A,E,U)$, where $V=\{(O_{j},M_{k})\;|\;O_{j}\in\mathcal{O},M_{k}\in\mathcal{M}_{j}\}$ represents the set of all operation-machine pairs, $A=\{(O_{l_{ij}},M_{k})\rightarrow(O_{i(j+1)},M_{k^{\prime}})\;|\;O_{l_{ij}},O_{i(j+1)}\in\mathcal{O};M_{k}\in\mathcal{M}_{l_{ij}};M_{k^{\prime}}\in\mathcal{M}_{l_{i(j+1)}}\}$ denotes the directed edge set, $E=\{(O_{j},M_{k})\leftrightarrow(O_{j^{\prime}},M_{k})\;|\;O_{j},O_{j^{\prime}}\in\mathcal{O};M_{k}\in\mathcal{M}_{j}\cap\mathcal{M}_{j^{\prime}}\}$ denotes the disjunctive (undirected) edge set, and $U=\{(O_{j},M_{k})\leftrightarrow(O_{j},M_{k^{\prime}})\;|\;O_{j}\in\mathcal{O},;M_{k},M_{k^{\prime}}\in\mathcal{M}_{j}\}$ connects all operation-machine nodes belonging to the same operation.
@@ Appendix B Neural Model for Scheduling Problems
For scheduling problems, we design an efficient neural model named MGL, which combines a multi-layer graph attention network (GAT) Veličković et al. (2018) encoder with a long short-term memory (LSTM) based Hochreiter & Schmidhuber (1997) context attention decoder.
@@ B.3 Comparison with Neural Model Used in SLIM
Our MGL is compared with the neural model used in SLIM Corsini et al. (2024), denoted as GAT-MHA. The main distinction between these models lies in their decoders, significantly impacting memory consumption and computational efficiency, as analyzed in Table 11. All data are collected from a $15\times 15$ instance, with the solution size $B$ set to 256.
@@ C.1 Formulations of Loss Functions
In SLIM Corsini et al. (2024), the locally optimal solution $y_{o}$ is treated as a pseudo-label, and the model is trained using cross-entropy loss. The loss function of SLIM can be expressed as::
@@
DPO Rafailov et al. (2024) employs a reference model $\pi_{\text{ref}}$, analogous to RLHF Stiennon et al. (2020), to regularize the trained model against excessive deviation from the initial policy. The DPO loss function is defined as:
@@
To simplify the training phase and align the training goal with the generation goal, SimPO Meng et al. (2024) eliminates the reference model and simplifies the loss function as follows:
@@
where $\beta$ is a constant that controls the scaling of the difference, and $\gamma$ is a target margin term. In contrast, POCO uses an adaptive objective gap factor to scale the differences, instead of relying on additional hyperparameters.
@@ C.2 Gradient Analysis
The gradient of $\mathcal{L}_{\text{POCO}}$ with respect to $\theta$ is:
$\nabla_{\theta}\mathcal{L}_{\text{POCO}}=\frac{\partial\mathcal{L}_{\text{POCO}}}{\partial z}\cdot\nabla_{\theta}z.$
@@
$\frac{\partial\mathcal{L}_{\text{POCO}}}{\partial z}=-(1-\sigma(z)).$
@@
$\nabla_{\theta}\mathcal{L}_{\text{POCO}}=\underbrace{-\frac{g(\bm{y}_{l})}{g(\bm{y}_{w})}}_{\text{Adaptive Scaling}}\cdot\underbrace{(1-\sigma(z))}_{\text{Confidence Weight}}\cdot\left(\underbrace{\frac{1}{|\bm{y}_{l}|}\nabla_{\theta}\log\pi_{\theta}(\bm{y}_{l}|\bm{x})-\frac{1}{|\bm{y}_{w}|}\nabla_{\theta}\log\pi_{\theta}(\bm{y}_{w}|\bm{x})}_{\text{Direction of Policy Update}}\right).$
@@
represent the confidence weight in SimPO and DPO.
DPO and SimPO rely on fixed hyperparameters (i.e., $\beta$ and $\gamma$) to control the gradient magnitude, requiring manual tuning and lacking dynamic adaptability. Furthermore, DPO does not perform length normalization on the log-likelihood calculations, which may lead to instability in scenarios with variable-length outputs. In contrast, POCO enhances robustness across diverse scenarios through adaptive scaling factors and length normalization, thereby reducing dependence on the $\beta$ hyperparameter.
@@ Appendix D Runtime Analysis for JSP
Shape OR-Tools L2S500 TGA500 L2S5k PDRs L2D SchN CL SLIM POCO SLIM POCO

### v3

@@ A.3 Flexible Job-shop Scheduling Problem
Disjunctive Graph for FJSP. To simplify the disjunctive graph for FJSP, we introduce operation-machine nodes, where an operation is decomposed into multiple operation-machine nodes. Each operation-machine node is similar to a node in JSP and can be treated as an action. Specifically, the disjunctive graph is formalized as $G=(V,A,E,U)$, where $V=\{(O_{j},M_{k})\;|\;O_{j}\in\mathcal{O},M_{k}\in\mathcal{M}_{j}\}$ represents the set of all operation-machine pairs, $A=\{(O_{l_{ij}},M_{k})\rightarrow(O_{l_{i(j+1)}},M_{k^{\prime}})\;|\;O_{l_{ij}},O_{l_{i(j+1)}}\in\mathcal{O};M_{k}\in\mathcal{M}_{l_{ij}};M_{k^{\prime}}\in\mathcal{M}_{l_{i(j+1)}}\}$ denotes the directed edge set, $E=\{(O_{j},M_{k})\leftrightarrow(O_{j^{\prime}},M_{k})\;|\;O_{j},O_{j^{\prime}}\in\mathcal{O};M_{k}\in\mathcal{M}_{j}\cap\mathcal{M}_{j^{\prime}}\}$ denotes the disjunctive (undirected) edge set, and $U=\{(O_{j},M_{k})\leftrightarrow(O_{j},M_{k^{\prime}})\;|\;O_{j}\in\mathcal{O},;M_{k},M_{k^{\prime}}\in\mathcal{M}_{j}\}$ connects all operation-machine nodes belonging to the same operation.
@@ Appendix B Neural Model for Scheduling Problems
For scheduling problems, we design an efficient neural model named MGL, which combines a multi-layer graph attention network (GAT) (Veličković et al., 2018) encoder with a long short-term memory (LSTM) based (Hochreiter & Schmidhuber, 1997) context attention decoder.
@@ B.3 Comparison with Neural Model Used in SLIM
Our MGL is compared with the neural model used in SLIM (Corsini et al., 2024), denoted as GAT-MHA. The main distinction between these models lies in their decoders, significantly impacting memory consumption and computational efficiency, as analyzed in Table 11. All data are collected from a $15\times 15$ instance, with the solution size $B$ set to 256.
@@ C.1 Formulations of Loss Functions
In SLIM (Corsini et al., 2024), the locally optimal solution $y_{o}$ is treated as a pseudo-label, and the model is trained using cross-entropy loss. The loss function of SLIM can be expressed as::
@@
DPO (Rafailov et al., 2023) employs a reference model $\pi_{\text{ref}}$, analogous to RLHF (Stiennon et al., 2020), to regularize the trained model against excessive deviation from the initial policy. The DPO loss function is defined as:
@@
To simplify the training phase and align the training goal with the generation goal, SimPO (Meng et al., 2024) eliminates the reference model and simplifies the loss function as follows:
@@
where $\beta$ is a constant that controls the scaling of the difference, and $\gamma$ is a target margin term. In contrast, BOPO uses an adaptive objective gap factor to scale the differences, instead of relying on additional hyperparameters.
Considering the similarity between the BOPO loss and that of policy gradients, we additionally analyze the loss function in the REINFORCE algorithm (Kwon et al., 2020) here:
$\mathcal{L}_{PG}(\pi_{\bm{\theta}}|\bm{x},\bm{y})=-(g(\bm{y})-b)\log\pi_{\bm{\theta}}(\bm{y}_{w}|\bm{x}),$ (8)
where $b$ is a baseline to distinguish positive or negative optimization signals for each sample $\bm{y}$. In POMO, $b=\sum_{i}^{B}g(\bm{y}_{i})/B$ is the average objective of $B$ samples. Note that, both BOPO and REINFORCE use the exact reward (objective) to optimal policy. However, BOPO directly distinguishes optimization signals based on exact preferences, and the strength of the optimization signal correlates with the difference in log-likelihood, requiring the model to maximize the probabilities gap between $\bm{y}_{w}$ and $\bm{y}_{l}$.
@@ C.2 Gradient Analysis
The gradient of $\mathcal{L}_{\text{BOPO}}$ with respect to $\theta$ is:
$\nabla_{\theta}\mathcal{L}_{\text{BOPO}}=\frac{\partial\mathcal{L}_{\text{BOPO}}}{\partial z}\cdot\nabla_{\theta}z.$
@@
$\frac{\partial\mathcal{L}_{\text{BOPO}}}{\partial z}=-(1-\sigma(z)).$
@@
$\nabla_{\theta}\mathcal{L}_{\text{BOPO}}=\underbrace{-\frac{g(\bm{y}_{l})}{g(\bm{y}_{w})}}_{\text{Adaptive Scaling}}\cdot\underbrace{(1-\sigma(z))}_{\text{Confidence Weight}}\cdot\left(\underbrace{\frac{1}{|\bm{y}_{l}|}\nabla_{\theta}\log\pi_{\theta}(\bm{y}_{l}|\bm{x})-\frac{1}{|\bm{y}_{w}|}\nabla_{\theta}\log\pi_{\theta}(\bm{y}_{w}|\bm{x})}_{\text{Direction of Policy Update}}\right).$
@@
represent the confidence weight in DPO and SimPO.
DPO and SimPO rely on fixed hyperparameters (i.e., $\beta$ and $\gamma$) to control the gradient magnitude, requiring manual tuning and lacking dynamic adaptability. Furthermore, DPO does not perform length normalization on the log-likelihood calculations, which may lead to instability in scenarios with variable-length outputs. In contrast, BOPO enhances robustness across diverse scenarios through adaptive scaling factors and length normalization, thereby reducing dependence on the $\beta$ hyperparameter.
@@ Appendix D Runtime Analysis for JSP
Shape OR-Tools L2S500 TGA500 L2S5k PDRs L2D SchN CL SLIM BOPO SLIM BOPO
@@ Appendix E Further Analysis of Hybrid Rollout
Appendix F Additional Experiments.
To further investigate the impact of $B$ on different problem types, we evaluate our method on TSP20/50 with $B=20/50$ (denoted as BOPO-). As shown in Table 13, when sampling the same number of solutions as POMO, the performance of the model remains comparable to $B=128$. This suggests that for small-scale or simple problems, the optimal $B$ is low, as the model can efficiently sample high-quality solutions with fewer times.
BOPO’s advantage primarily stems from pairwise preference learning, rather than simply being an RL variant with the proposed filtering method. To demonstrate this point, we compare BOPO with a POMO variant employing our Hybrid Rollout and Uniform Filtering (denoted as POMO+). The key difference lies in the loss: while BOPO uses pairwise loss, POMO+ calculates a standard POMO loss based on the filtered $K$ solutions, as it lacks the pairwise mechanism. Results in Table 13 show that POMO+ is inferior to BOPO, demonstrating the effectiveness of the preference learning of BOPO. Moreover, POMO+ is even inferior to POMO, underscoring that the filtering method is tightly adapted to BOPO and may not be applicable to general RL methods.
Table 13: Average gaps (%) on generated TSP instances.
Methods TSP20 TSP50
POMO 0.002 0.042
POMO+ 0.005 0.081
BOPO- 0.000 0.009
BOPO 0.000 0.009